You're using a convex lens with f = 10.0 cm to stu
Answer #1

Image being upright for a convex lens implies that it is a virtual image.

f = 10 cm

Since the magnification M = v/u=-1.9

v = -1.9\times u

where f is the focal length, v is the image distance and u is the objecct distance. Substituing in the lens formula,

\frac{1}{u}+\frac{1}{v}=\frac{1}{f}

Substituing and solving, we get

u=4.74cm